Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for August 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, August 26, 1998

FOR DATES: From 7:00 a.m. Wednesday, August 19, 1998 to 7:00 a.m. Wednesday, August 26, 1998.

The Thermal Destruction Facility (TDF) was on-line for 120.81 hours and processed 4780.7 tons of contaminated soil during the time frame of this report. The TDF was on-line 71.9 % of the time with an average hourly feed rate of 39.6 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 111,588.7 tons of contaminated soil, which amounts to 40.4 % of the overall contaminated soil on-site.

As indicated on the previous week's report, the contractor stopped feed on Wednesday, August 19, 1998 in preparation to perform maintenance on the TDF. The maintenance down time was primarily attributed to replacing the worn feed screws. The feed screws were replaced with a previously used pair of feed screws that was modified. Some of the modifications to the screws included the welding on new flights and applying a hard face metal to the tips of the flights. There was a delay in the installation of the new feed screws, as additional machining of the interior of the screw's shaft was required for a proper fit. Preventive maintenance was also performed on the Induced Draft (ID) fan motor, the feed screw drive chain and other TDF system components.

During the down time, the contractor performed the monthly calibrations on the TDF monitoring equipment and performed the monthly interlock check. On Wednesday, August 19, at 1527 hours, the contractor initiated a planned Thermal Relief Valve (TRV) opening as part of the monthly interlock checks. The appropriate authorities were notified and feed to the incinerator had been shut down for approximately 5 hours prior to the opening. Hazardous waste feed to the incinerator resumed at 0434 hours on Friday.

Additional feed stoppages during the time frame of this report are attributed to the manual stoppage of feed as a precaution during local lightning storm activity and to perform minor preventive maintenance to the TDF equipment.

In response to detecting volatile organic readings greater than 1 ppm at the Perimeter Air Monitoring (PAM) stations, Summa canisters were deployed to collect grab samples of the air upwind and downwind from the site. In all instances, the speciated volatile organic compound readings were well below any action level of concern. In the near future, the results of the canister samples will be available and will be provided to ATSDR.

Construction activities during the time frame of this report included wastewater management, excavation within the exclusion zone, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, rock and concrete debris handling, cleaning and backfilling. Concover, an odor and dust suppressant was placed on the ash stockpile to prevent fugitive emissions and soil erosion. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

No drums were excavated during the time frame of this report. Demolition debris continued to be unearthed and separated in the exclusion zone. Two rolloffs of PPE and miscellaneous debris were transported off-site to a hazardous waste landfill.

The weekly progress meeting was held on Tuesday, August 25, 1998 between representatives from the USEPA, USACE, PADEP, and OHM.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE


Thursday, August 20, 1998

FOR DATES: From 7:00 a.m. Wednesday, August 12, 1998 to 7:00 a.m. Wednesday, August 19, 1998.

The Thermal Destruction Facility (TDF) was on-line for 155.5 hours and processed 5821.5 tons of contaminated soil during the time frame of this report. The TDF was on-line 92.6 % of the time with an average hourly feed rate of 37.4 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 107,193 tons of contaminated soil, which amounts to 39% of the overall contaminated soil on-site.

In response to a number of feed screw stoppages during and prior to the time frame of this report, OHM's engineers and technical staff reviewed the feed screw design. It was decided to replace the current feed screws with a previously used pair of feed screws that had been modified. Some of the modifications to the screws included the welding of new flights onto the previously used feed screws and applying a hard face metal to the tips of the flights.

A planned shutdown occurred Wednesday afternoon in order to replace the feed screws and perform maintenance to the TDF. It is expected that waste feed will resume on Friday, August 21, or Saturday, August 22, 1998.

Construction activities during the time frame of this report included wastewater management, the restarting of mass excavation within the exclusion zone, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, rock and concrete debris handling, cleaning and backfilling. Concover, an odor and dust suppressant, was placed on the ash stockpile to prevent fugitive emissions and soil erosion. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

As noted on the Situation Report for Saturday, August 15, 1998, grab type air samples (using Summa canisters) have been collected at some perimeter air monitoring stations. It should be noted that these grab samples have been collected in response to ATSDR comments after their review of air monitoring plans and data reports for the Drake project. The Saturday Situation Report is further clarified by noting that the Summa canisters were not required to be collected for regulatory purposes. Analytical results should be available for a number of such sampling events in a couple weeks. Results will be forwarded to ATSDR.

No drums were excavated during the time frame of this report. Demolition debris continued to be unearthed in the exclusion zone. The demolition debris includes ductile iron water pipe, train rails, and miscellaneous building debris.

Related site activities during this period included:

* A weekly progress meeting held on Tuesday, August 18, 1998.

* A field management partnering meeting was held on Wednesday, August 12, 1998. Represented were the USEPA, USACE, PADEP, and OHM.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE


Thursday, Auguat 13, 1998

FOR DATES: From 7:00 a.m. Wednesday, August 5, 1998 to 7:00 a.m. Wednesday, August 12, 1998.

The Thermal Destruction Facility (TDF) was on-line for 162.8 hours and processed 6485 tons of contaminated soil during the time frame of this report. The TDF was on-line 96.9 % of the time with an average hourly feed rate of 39.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 102,498 tons of contaminated soil, which amounts to 37% of the overall contaminated soil on-site.

In response to a number of feed screw stoppages, the contractor is reviewing soil feed conditions and feed screw operation, maintenance and design parameters.

Construction activities during the time frame of this report included wastewater management, preparations in the exclusion zone to continue excavation on the south face of the soil stockpile, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, rock and concrete debris handling, cleaning and backfilling. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

No drums were excavated during the time frame of this report. A large amount of demolition debris was unearthed in the exclusion zone. The demolition debris includes ductile iron water pipe, train rails, and miscellaneous debris that was unearthed during earlier excavation activities by the contractor and then buried in the stockpile.

Related site activities during this period included a weekly progress meeting held on Tuesday, August 11, 1998, and a field management partnering meeting on Wednesday, August 12, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, August 5, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 29, 1998 to 7:00 a.m. Wednesday, August 5, 1998.

The Thermal Destruction Facility (TDF) was on-line for 63.14 hours and processed 2444.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 37.6 % of the time with an average hourly feed rate of 38.7 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 96,004.7 tons of contaminated soil, which amounts to approximately 34.8 % of the overall contaminated soil on-site.

For the beginning of this reporting period, the TDF remained in a cold shutdown mode in order to complete maintenance to the TDF. The maintenance included the replacement of the kiln feed screws, the replacement of the kiln feed screw cooling jacket, maintenance to the Induced Draft (ID) fan motor, and preventative maintenance on many other components of the TDF. On Friday at 1705 hours, heat up of the TDF began. Hazardous waste feed to the TDF was restarted at 0952 hours on Sunday, August 2, 1998.

Since operations restarted, the TDF has been on-line 90.5 % of the time. Off-line time was necessary to perform repairs to the weigh belt speed sensor.

Construction activities during the time frame of this report included wastewater management, minor excavation of soils within the exclusion zone, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, and rock and concrete debris handling, cleaning and backfilling. The contractor continued to perform real time perimeter air monitoring and off-site air monitoring activities.

Related site activities during this period included a weekly progress meeting held on Tuesday, August 4, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.